mathslover (mathslover):

The radius of the first circle is 1 cm, that of the second is 1/2 cm and that of the third is 1/4 cm and so on indefinitely. The sum of the areas of the circles is ?

Parth (parthkohli):

\[\pi + {\pi \over 4} + {\pi \over 16}\cdots\]

Parth (parthkohli):

Hmm...

mathslover (mathslover):

You mean to say like this : \[\large{S = \pi + \frac{\pi }{4} + \frac{\pi}{16} + .... }\]

mathslover (mathslover):

But what next?

Parth (parthkohli):

I think it's approaching something.

mathslover (mathslover):

Well i think we can take : infinite geometric progression ...

mathslover (mathslover):

pi ( 1 + 1/4 + 1/16 + .... )

Parth (parthkohli):

Where did you get this question?

mathslover (mathslover):

I have a book, Math IQ Challenge

Parth (parthkohli):

It'd be \(\dfrac{4}{3}\pi\)

Parth (parthkohli):

lol - we solved it before asking it on M.SE :)

mathslover (mathslover):

:) \[\large{ S _\infty = \frac{a}{1-r}}\] -- infinite geometric series a = 1 r = 1/4 1 / ( 1- 1/4 ) = 1/ (3/4) = 4/3

mathslover (mathslover):

Yeah ... thanks @ParthKohli